A laboratory technician needs to make a 32-liter batch of a 40% acid solution. How can the laboratory technician combine a batch of an acid solution that is pure acid with another that is 20% to get the desired concentration?
Answer by richwmiller(17219)

a+b=32,
100*a+20*b=40*32
a=32-b
100*(32-b)+20*b=1280
3200-100b+20*b=1280
-80*b=-1920
b=24
a=32-b
a=8 liter at 100%
b=24 liter at 20%
check
100*8+20*24=40*32
800+480=1280
1280=1280
ok